# SentenceVAE: Enable Next-sentence Prediction for Large Language Models with Faster Speed, Higher Accuracy and Longer Context

### v3

@@ Abstract
Current large language models (LLMs) primarily utilize next-token prediction method for inference, which significantly impedes their processing speed. In this paper, we introduce a novel inference methodology termed next-sentence prediction, aimed at enhancing the inference efficiency of LLMs. We present Sentence Variational Autoencoder (SentenceVAE), a tiny model consisting of a Sentence Encoder and a Sentence Decoder. The Sentence Encoder can effectively condense the information within a sentence into a singular token, while the Sentence Decoder can reconstruct this compressed token back into sentence. By integrating SentenceVAE into the input and output layers of LLMs, we develop Sentence-level LLMs (SLLMs) that employ a sentence-by-sentence inference method. In addition, the SentenceVAE module of SLLMs can maintain the integrity of the original semantic content by segmenting the context into sentences, thereby improving accuracy while boosting inference speed. Moreover, compared to previous LLMs, SLLMs process fewer tokens over equivalent context length, significantly reducing memory demands for self-attention computation and facilitating the handling of longer context. Extensive experiments on Wanjuan dataset have reveal that the proposed method can accelerate inference speed by 204$\sim$365%, reduce perplexity (PPL) to 46$\sim$75of its original metric, and decrease memory overhead by 86$\sim$91for the equivalent context length, compared to the token-by-token method.
@@ Introduction
We revisited the mechanisms underlying human interaction. Typically, when individuals prepare to speak, they have already formulated the semantic content of an entire sentence in their minds. However, due to the limitations of the vocal cords, which are relatively inefficient organs for sound production, speakers are compelled to articulate sentences word-by-word, much like the token-by-token generation method of LLMs. Similarly, listeners tend to comprehend sentences holistically, rather than word-by-word. This observation suggests that sentences, rather than individual words, constitute the fundamental units of comprehension in natural language. Motivated by this insight, we developed the Sentence Variational Autoencoder (SentenceVAE), as depicted in Fig. 1. This model comprises a Sentence Encoder and a Sentence Decoder. The encoder compresses the multiple tokens of a sentence into a single sentence-level token, mirroring the processing capabilities of the human auditory system, while the decoder reconstructs the sentence’s semantic content from this token, analogous to the function of the vocal cords. By integrating the encoder and decoder components of SentenceVAE seamlessly into the input and output layers of LLMs, these enhanced models are capable of operating within a more efficient sentence-level embedding space, performing next-sentence prediction.
@@
  • Extensive experiments conducted on various LLMs have demonstrated that the proposed method significantly accelerates inference speed while preserving the accuracy of these models.
@@
In this section, we introduce the SentenceVAE framework. Embedding this framework into a LLM enables the model to implement a sentence-by-sentence prediction method, significantly accelerating its inference speed. The inference process of the proposed method compared to published LLMs is illustrated in Fig. 2. It is evident that for the same context, an LLM embedded with SentenceVAE requires only three inference steps to obtain the inference result, whereas previous LLMs necessitate twenty-three inference steps. Furthermore, the proposed method divides the context into sentences without compromising the original semantics, thereby enhancing the inference speed of LLMs while preserving or even enhancing accuracy.
@@
Our SentenceVAE primarily consists of a Sentence Encoder and a Sentence Decoder. The ecoder encodes multiple word-level tokens from a sentence into a single sentence-level token, while the decoder reconstructs this sentence-level token back into the original sequence of word-level tokens. To ensure that the multiple tokens input to the encoder come from individual sentences, we propose a specialized sentence segmentation mechanism. Additionally, we introduce a feature fusion mechanism, which enable the encoder to encode variable-length token sequences into a single sentence-level token.
@@
To derive the sentence embedding vector, we fuse these $L$ features into a single vector $\boldsymbol{\Omega}_{t}\in\mathbb{R}^{\text{hidden\_size}}$. This sentence embedding vector is then fed into the decoder-only LLM, which generates a new sentence embedding vector $\boldsymbol{\Omega}_{t+1}$ at each time step.
Feature Fusion Mechanism: after the encoder blocks generates $\boldsymbol{H}$, we propose a method to fuse them into a single sentence embedding vector $\boldsymbol{\Omega}_{t}$. This method involves accumulating the $L$ vectors and normalizing them using Layer Normalization (Ba, Kiros, and Hinton 2016), as shown in Eq. 3.
@@
In the prediction phase, given $\boldsymbol{\Omega}_{t+1}$, initialize the input token $d_{0}=<bos>$, the decoder outputs $d_{1}$, and then input $[d_{0},d_{1}]$ and $\boldsymbol{\Omega}_{t+1}$ into the decoder. Repeat this process until $d_{l^{\prime}+1}=<eos>$, and terminate the iteration, as shown in Eq.4.
@@
Termination judgment layer: in previous LLMs, the final fully connected layer $\textbf{FC}_{\text{llm}}$ typically converts an $\boldsymbol{H}_{\text{llm}}\in\mathbb{R}^{t\times\text{hidden\_size}}$ to a probability vector (logits) $\boldsymbol{P}_{\text{llm}}\in\mathbb{R}^{t\times V}$, where $V$ represents the size of the vocabulary. This allows the model to determine the next token output using sampling algorithms, e.g., Greedy Search, Beam Search (Graves 2012), Random Sampling, Temperature Sampling (Hinton, Vinyals, and Dean 2015), Top-K Sampling (Fan, Lewis, and Dauphin 2018), and Nucleus Sampling (Holtzman et al. 2019). If the current token output is a special token like eos (end-of-sequence), the iteration terminates. Notably, in SLLMs, the output of $\textbf{DB}_{\text{llm}}$ is a sentence-level hidden state vector $\boldsymbol{H}_{\text{sllm}}$. Therefore, traditional token generation methods are not applicable. Instead, we use a new fully connected layer $\textbf{FC}_{\text{sllm}}$ called termination judgment layer that converts the $\boldsymbol{H}_{\text{sllm}}$ into a 2-dimensional boolean vector $\boldsymbol{B}_{\text{stop\_flag}}$, as shown in Eq. 8. This vector helps determine whether the current sentence-level hidden state vector signifies the end of a sentence (stop flag) or needs further decoding by the sentence decoder. If the vector indicates an end flag, the iteration terminates. Otherwise, the sentence decoder processes the embedding to generate corresponding tokens, as shown in Eq. 9.
@@
A series of word-level tokens can be represented by a sentence-level token with strong robustness
@@
Table 2: Output examples of SentenceVAEs.
@@ 1 Hello, 3 Hello, Hello, Hello,
The evaluation results affirm that SentenceVAEs achieve their intended functionality, substantiating their theoretical underpinnings. Table 2 showcases a variety of test cases, including Samples 1, 2, 3, 5, 6, which serve as conventional examples. To evaluate model robustness, several out-of-distribution samples, marked in blue, are also incorporated:
@@
SLLMs can work in setence-level embedding space with faster inference speed, more accurate PPL, and longer context
@@
Longer Context: the maximum context length manageable by a model is generally limited by GPU memory availability. The SLLM framework compresses multiple original tokens into a single token, reducing memory usage for equivalent-length contexts. This compression allows for an extended context capacity within the same hardware constraints.
SLLMs still follow the Scaling Law
During the training of SLLMs, an analysis of the loss curve indicates compliance with the Scaling Law, as depicted in Figure 3. This observation supports the hypothesis that the SLLM framework is amenable to effective scaling up to larger language models.
@@
This section encapsulates the contributions of our research and delineates several promising avenues for future exploration.

### v5

@@ Abstract
Current large language models (LLMs) primarily utilize next-token prediction method for inference, which significantly impedes their processing speed. In this paper, we introduce a novel inference methodology termed next-sentence prediction, aiming at enhancing the inference efficiency of LLMs. We present Sentence Variational Autoencoder (SentenceVAE), which includes a Sentence Encoder to compress multiple tokens in a sentence into a single token, and a Sentence Decoder to reconstruct it. By integrating SentenceVAE into the input and output layers of LLMs, we develop Sentence-level LLMs (SLLMs) that employ a sentence-by-sentence inference method. In addition, the SentenceVAE module of SLLMs can maintain the integrity of the original semantic content by segmenting the context into sentences, thereby improving accuracy while boosting inference speed. Moreover, compared to previous LLMs, SLLMs process fewer tokens over equivalent context length, significantly reducing memory demands for self-attention computation and facilitating the handling of longer context. Extensive experiments on Wanjuan dataset have revealed that the proposed method can accelerate inference speed by 204$\sim$365%, reduce perplexity (PPL) to 46$\sim$75of its original metric, and decrease memory overhead by 86$\sim$91for the equivalent context length, compared to previous token-by-token methods.
@@ Introduction
To enhance inference speed while preserving or even enhancing accuracy, we explore method for adaptively selecting the optimal number of predicted tokens based on the semantic content of the text. Consequently, we introduce Sentence Variational Autoencoder (SentenceVAE), which comprises a sentence encoder and a sentence decoder. The schematic of SentenceVAE is illustrated in Fig. 1. The sentence encoder is designed to condense the information of an entire sentence into a single token, while the sentence decoder reconstructs this compressed token back into a sentence. By seamlessly incorporating the encoder and decoder components of SentenceVAE into the input and output layers of LLMs, these enhanced models are capable of performing next-sentence prediction. This integration not only enriches the models’ capabilities but also accelerates the inference speed, leading to improved efficiency in language understanding and generation. Moreover, since SentenceVAE segments text at the sentence level, the semantic integrity is better preserved, ensuring the accuracy of the inferrence. Specifically, prior to inputting text into LLMs, SentenceVAE segments the text into sentences and utilizes the sentence encoder to compress the information contained within each sentence into a single token. These compressed tokens are then fed as input to the LLMs. The LLMs predict their output based on these compressed tokens. Finally, the sentence decoder decodes the predicted token into the final output of the LLMs.
@@
  • Extensive experiments conducted on various LLMs have demonstrated that the proposed method not only accelerates inference speed, but also improves accuracy, and reduces memory usage, enabling the model to handle longer contexts with the same device memory.
@@
We revisited the mechanisms underlying human interaction. Typically, when individuals prepare to speak, they have already formulated the semantic content of an entire sentence in their minds. However, due to the limitations of the vocal cords, which are relatively inefficient organs for sound production, speakers are compelled to articulate sentences word-by-word, much like the token-by-token generation method of LLMs. Similarly, listeners tend to comprehend sentences holistically, rather than word-by-word. This suggests that sentences are the fundamental units of comprehension in natural language. Inspired by this, we developed the Sentence Variational Autoencoder (SentenceVAE), which includes a Sentence Encoder to compress a sentence into a single token and a Sentence Decoder to reconstruct it. By integrating SentenceVAE into LLMs, these models can operate within a more efficient sentence-level embedding space, enabling next-sentence prediction. This method significantly reduces the number of tokens for the same context length, thereby improving the model’s inference speed and reducing the memory usage required for self-attention calculation, enabling the model to handle longer contexts. Additionally, by understanding language at the sentence level, the model demonstrates improved accuracy.
@@
Our SentenceVAE primarily consists of a Sentence Encoder and a Sentence Decoder. The encoder encodes multiple word-level tokens from a sentence into a single sentence-level token, while the decoder reconstructs this sentence-level token back into the original sequence of word-level tokens. To ensure that the multiple tokens input to the encoder come from individual sentences, we propose a specialized sentence segmentation mechanism. Additionally, we introduce a feature fusion mechanism, which enable the encoder to encode variable-length token sequences into a single sentence-level token.
@@
To derive the sentence-level token, we fuse these $L$ features into a single vector $\boldsymbol{\Omega}_{t}\in\mathbb{R}^{\text{hidden\_size}}$. This sentence-level token is then fed into the decoder-only LLM, which generates a new sentence-level token $\boldsymbol{\Omega}_{t+1}$ at each time step.
Feature Fusion Mechanism: after the encoder blocks generates $\boldsymbol{H}$, we propose a method to fuse them into a single sentence-level token $\boldsymbol{\Omega}_{t}$. This method involves accumulating the $L$ vectors and normalizing them using Layer Normalization (Ba, Kiros, and Hinton 2016), as Eq. 3.
@@
In the prediction phase, given $\boldsymbol{\Omega}_{t+1}$, initialize the input token $d_{0}=<bos>$, the decoder outputs $d_{1}$, and then input $[d_{0},d_{1}]$ and $\boldsymbol{\Omega}_{t+1}$ into the decoder. Repeat this process until $d_{l^{\prime}+1}=<eos>$, and terminate the iteration, as Eq. 4.
@@
Termination judgment layer: in previous LLMs, the final fully connected layer $\textbf{FC}_{\text{llm}}$ typically converts an $\boldsymbol{H}_{\text{llm}}\in\mathbb{R}^{t\times\text{hidden\_size}}$ to a probability vector (logits) $\boldsymbol{P}_{\text{llm}}\in\mathbb{R}^{t\times V}$, where $V$ represents the size of the vocabulary. This allows the model to determine the next token output using sampling algorithms, e.g., Greedy Search, Beam Search (Graves 2012), Random Sampling, Temperature Sampling (Hinton, Vinyals, and Dean 2015), Top-K Sampling (Fan, Lewis, and Dauphin 2018), and Nucleus Sampling (Holtzman et al. 2019). If the current token output is a special token like eos (end-of-sequence), the iteration terminates. Notably, in SLLMs, the output of $\textbf{DB}_{\text{llm}}$ is a sentence-level hidden state vector $\boldsymbol{H}_{\text{sllm}}$. Therefore, traditional token generation methods are not applicable. Instead, we use a new fully connected layer $\textbf{FC}_{\text{sllm}}$ called termination judgment layer that converts the $\boldsymbol{H}_{\text{sllm}}$ into a 2-dimensional boolean vector $\boldsymbol{B}_{\text{stop\_flag}}$, as shown in Eq. 8. This vector helps determine whether the current sentence-level hidden state vector signifies the end of a sentence (stop flag) or needs further decoding by the sentence decoder. If the vector indicates an end flag, the iteration terminates. Otherwise, the sentence decoder processes the embedding to generate corresponding tokens, as Eq. 9.
@@
Sentence-level Tokens
@@
Table 2: Test examples of SentenceVAEs. Text in blue indicates out-of-distribution samples, text in red highlights partial mismatches between output and input, and the denotes missing output.
@@ 1 Hello, 3 Hello, Hello, Hello,
The evaluation results confirm that SentenceVAEs function as intended,validating their theoretical foundations. Table 2 showcases a variety of test cases, including Samples 1, 2, 3, 5, 6, which serve as conventional examples. To evaluate model robustness, several out-of-distribution samples, marked in blue, are also incorporated:
@@
Overall, the experimental results indicate that, a series of word-level tokens can be represented by a sentence-level token with strong robustness.
Sentence-level LLMs (SLLMs)
@@
Longer Context: The maximum context length a model can handle is typically constrained by GPU memory availability. The SLLM framework addresses this by compressing multiple original tokens into a single token, reducing memory usage for contexts of equivalent length. This compression enables an extended context capacity within the same hardware limitations.
In summary, SLLMs can work in setence-level embedding space with faster inference speed, more accurate PPL, and longer context.
Scaling Law of SLLMs
During the training of SLLMs, an analysis of the loss curve indicates compliance with the Scaling Law, as depicted in Figure 3. This result indicates that SLLMs also adhere to the Scaling Law, suggesting that this theory can be extended to larger models.
@@
This section encapsulates the contributions of our research and delineates several promising avenues for exploration.